The firewall has remained in a stable state and changes are being completed as normal.
Nov 24, 10:03 GMT
The change freeze has now been lifted into a controlled state. Our engineers continue to closely monitor the environment.
Nov 2, 15:48 GMT
Further performance issues have been seen on our core firewalls. We have now reapplied the change freeze. We continue to work closely with partners to resolve the issue.
Oct 31, 13:31 GMT
Our core firewalls have remained stable with no impact on performance. The change freeze has now been lifted and change requests are being processed as normal for Luton, West Berks and our core customers set.
Oct 27, 20:00 GMT
We are still experiencing a software issue on our core firewalls. While the issue persists, we are unable to process any firewall change requests for Luton, West Berks and customers with connectivity directly through RM. Our infrastructure development team are actively engaged with Cisco to resolve the issue.
Oct 24, 11:29 GMT
We will be implementing maintenance on a core firewall at 8pm tonight to resolve the issue. We are anticipating downtime up to 20 minutes during this change.
Oct 21, 12:23 GMT
We have identified a possible software issue in one of the firewalls which is preventing us from adding new changes to it. During this period the network and access will remain fully operational. Our Engineers are currently working with 3rd party vendors and Cisco to fix the issue identified.
Oct 21, 09:27 GMT